OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through the given points. 1. (1,3) (7,4) 2. (0,-3) (-5,0) 3. (-6,-7) (-5,1) 4. (4,2) (7, -4)

OpenStudy (♪chibiterasu):

There is no given slope?

OpenStudy (anonymous):

Nope.

OpenStudy (♪chibiterasu):

Alright, then m remains the same.

OpenStudy (anonymous):

What does that mean? I'm sorry, I'm horrible at math.

OpenStudy (♪chibiterasu):

slope intercept is y = mx + b y in the y-int m is the slope x is the x-int and I forgot what b is, but thats not important.

OpenStudy (anonymous):

Okay, now what?

OpenStudy (♪chibiterasu):

First we will convert it into point-slope form which will get us to its slope-intercept form

OpenStudy (♪chibiterasu):

(y - y^1) = m(x - x^1)

OpenStudy (♪chibiterasu):

(1,3) (7,4) (3 - 4) = m(1 - 7) -1 = 1m - 7

OpenStudy (♪chibiterasu):

That is the solution for 1

OpenStudy (♪chibiterasu):

Do you think you can try to solve 2?

OpenStudy (anonymous):

But, it says slope-intercept form. So isn't it supposed to be in y= format?

OpenStudy (♪chibiterasu):

we do point-slope first and then simplify it. Once simplified it becomes slope-intercept

OpenStudy (texaschic101):

you have to find the slope...do you know how to do that ?

OpenStudy (texaschic101):

to find the slope.use this formula : slope(m) = (y2 - y1) / (x2 - x1) (1,3) -- x1 = 1 and y1 = 3 (7,4) -- x2 = 7 and y2 = 4 now lets sub it into the formula and find the slope slope(m) = (4 - 3) / (7 - 1) m = 1/6 <-- this is your slope now we will use this formula : y - y1 = m(x - x1) slope(m) = 1/6 and you can use wither of your points, I will use (1,3) (1,3) -- x1 = 1 and y1 = 3 now lets sub y - 3 = 1/6(x - 1) y - 3 = 1/6x - 1/6 y = 1/6x - 1/6 + 3 y = 1/6x - 1/6 + 18/6 y = 1/6x + 17/6 <-- slope intercept form

OpenStudy (texaschic101):

you have to do that for each problem